# arXiv:2603.01543v1[math.DG] 02 Mar 2026

### v1

@@ 1.5. The $1$-harmonic Mass
with $V\in\mathscr{C}^{\infty}(M)$. If $\Lambda>0$, $M$ is compact and $\partial M$ is connected, Boucher, Gibbons, and Horowitz have established the following rigid upper bound
$|\partial M|\,\leq\,\frac{12\pi}{\Lambda}\,.$
It follows that:
Theorem 1.6.
The only compact three-dimensional static triple with $\Lambda>0$ and smooth connected boundary for which the $1$-harmonic Mass is vanishing at some point is the de Sitter solution, i.e. the round hemisphere with constant sectional curvatures equal to $\Lambda/3$.
@@
In Section 2, we establish monotonicity formulas on Riemannian bands under the assumption of a lower bound on the scalar curvature. These formulas are substantially more general than those required for the arguments developed in the rest of the paper. Nonetheless, they serve a dual purpose: on the one hand, they situate our contribution within a broader research perspective, and on the other hand, they are expected to play a central role in forthcoming extensions of the theory in several natural directions, including the analysis of Penrose-type inequalities in the presence of a nonzero cosmological constant. In Section 3, we specialize to the case of globally defined $p$-Green’s functions with $1<p<3$. Starting from the study of rotationally symmetric solutions, we then identify a suitable choice of the structural coefficients that appear in the general formulation of the monotonicity formulas, in such a way as to characterize static zero-mass metrics (space forms). These metrics constitute the natural reference configurations for any proposed notion of mass and are therefore required to arise as the rigidity case in inequalities intended to yield positive mass theorems. In Section 4, we begin our detailed analysis of the case of a positive cosmological constant, which is the main focus of the present work, and we refine the description of the structural coefficients. These will be crucial for evaluating the limits of the monotone quantities at both the pole of the Green’s function and the boundary of the manifold. This, in turn, prepares the ground for the proof of the main results, namely Theorems 1.3 and 1.4, corresponding, respectively, to Theorem 5.4 in Section 5 and Theorem 6.2 in Section 6. In Section 4, we also provide additional remarks on the behavior of the structural coefficients in the limit as $p\to 1^{+}$, which support some of the more speculative considerations presented in Subsection 1.4. Appendix A contains auxiliary material needed for the treatment of the case $\Lambda>0$ in Theorem 3.10, as well as for the analysis carried out in Section 4. The material in Appendix B underpins the arguments developed in Sections 5 and 6.
@@ 2. Monotonicity formulas for $p$-harmonic functions on Riemannian bands
where the $p$-Laplacian is defined as $\Delta_{p}\phi={\rm div}(|\nabla\phi|^{p-2}\nabla\phi)$. The fact that the above problem admits a unique solution $u\in\mathcal{W}^{1,p}(\Omega)\cap\mathscr{C}^{0}(\overline{\Omega})$ is standard, see for example by [49, Theorem 2.16] for a proof based on the direct method in the calculus of variations. Following [55], we set $w=-(p-1)\log u$, so that $w$ satisfies
@@ 3.2. Green’s functions for the $p$-Laplacian on 3-dimensional spaceforms and the selection of $\alpha$
one concludes that $\Psi(r)>0$ for all $0<r<R_{\Lambda}$. Having all these elements at our disposal, we can finally invert relations (3.36), thereby constructing two solutions $\lambda$ and $\mu$ of system (2.4), defined for all times and exhibiting the prescribed asymptotic expansions as $t\to$-$\infty$. This ensures that these globally defined solutions are indeed the maximal extensions fo the ones constructed in Theorem 3.4 and Theorem 3.7
@@ 5. A Positive Mass Theorem for the Polarized $p$-harmonic Mass
Let $(M,g)$ be a compact three-dimensional Riemannian manifold with smooth boundary $\partial M$. For $1<p<3$, the Polarized $p$-harmonic Mass of $(M,g)$, with pole at $x\in M\setminus\partial M$ is given by
@@
Let $(M,g)$ be a compact three-dimensional Riemannian manifold with compact smooth boundary $\partial M$, whose scalar curvature $\mathrm{R}$} satisfies
@@ 6. A Positive Mass Theorem for the $p$-harmonic Total Mass
Let $(M,g)$ be a compact three-dimensional Riemannian manifold with smooth minimal boundary $\partial M$, whose scalar curvature $\mathrm{R}$} satisfies
@@
Theorem B.2, coupled with the Comparison Principle for $p$-harmonic functions (see, e.g.  [60, Theorem 3.5.1]), implies that for every $n$ there exists $M_{n}$ such that
@@
Fix a point $x_{0}\in\partial M$, and let $\nu$ be the unit normal vector to $\partial M$ at $x_{0}$, pointing towards the interior of $M$. It is evident that establishing the desired statement is equivalent to proving that
@@
Now, the upper bound in (6.10) coupled with the Comparison Principle for $p$-harmonic functions (see, e.g.  [60, Theorem 3.5.1]) implies that the functions $v_{\varepsilon}$ are uniformly bounded in the complement of suitably small coordinate balls centered at $x_{\varepsilon}$, namely

### v2

@@ 1.5. The $1$-harmonic Mass
Figure 1. Graph of the function $p_{\Lambda}(r)=\frac{r}{2}\left(1-\frac{\Lambda}{3}r^{2}\right)$ over the interval $0,\sqrt{3/\Lambda}]$. The black dashed lines highlight the maximum point and the maximal value of the function; the blue dashed lines represent the level $m$ and its two positive preimages $R_{\Lambda}^{-}(m)$ and $R_{\Lambda}^{+}(m)$.
Having introduced the appropriate notation, we can now confirm the claimed consistency check for the 11‑harmonic mass of the Schwarzschild–de Sitter solutions. For this purpose, it is convenient to attach a spherical cap (hemisphere) of radius $R_{\Lambda}^{-}(m)$ to the black‑hole horizon of $g_{\Lambda,m}$ – that is, to the boundary component corresponding to $r=R_{\Lambda}^{-}(m)$ – thereby producing a manifold with connected boundary and a distinguished north pole. In this configuration, one verifies that the IMCF starting from the north pole $N$ of the spherical cap and terminating at the cosmological horizon $\{r=R_{\Lambda}^{+}(m)\}$ is still described by
$w_{N}^{(1)}=2\log r,$
as in the case $m=0$. The only modification concerns the maximal existence time of the flow, which is now given by
$T^{*}(N)=T_{\Lambda}^{+}(m)=2\log R_{\Lambda}^{+}(m).$
Moreover, for $\infty<\tau\leq T_{\Lambda}^{+}(m)$, the areas of the corresponding level sets remain equal to $4\pi e^{\tau}$. Consequently, performing in the present setting the very same integration that we performed in the massless case yields that the value of the $1$-harmonic Mass is now given by
$\frac{R^{+}_{\Lambda}(m)}{2}\left(1-\frac{\Lambda}{3}\bigl(R^{+}_{\Lambda}(m)\bigr)^{2}\right)=m\,$,
as wished. Beyond this consistency verification, we observe that the notion of 11‑harmonic mass admits a coherent and natural extension to the setting of three‑dimensional Riemannian bands $(M,g)$ with scalar curvature satisfying $\mathrm{R}_{g}\geq 2\Lambda>0$ and boundary
$\partial M=\partial M^{-}\sqcup\partial M^{+}\,.$
In this more general framework, we will demonstrate that the 11‑harmonic mass satisfies a Riemannian Penrose‑type inequality, in agreement with the conjectural inequalities proposed by Bray and Chruściel in [18] (see in particular their Penrose conjecture, formula (4.7)). The primary formal distinction between the definition given in (1.11), for the case of connected boundary, and the one we are about to introduce lies in the specification of the standard existence interval and the consequent normalization of the initial mass content. To clarify these concepts, we assume that both boundary components are smooth closed minimal surfaces and we suppose without loss of generality that
$|\partial M^{-}|\leq|\partial M^{+}|\,$,
thus interpreting the boundary component $\partial M^{-}$ as a black-hole–type horizon and the boundary component $\partial M^{+}$ as a cosmological horizon. Under this hypothesis, we define the real numbers $R^{-}\!=R^{-}(M,g)$ and $T^{-}\!=T^{-}(M,g)$ as
$R^{-}\!=\sqrt{\frac{|\partial M^{-}|}{4\pi}}\qquad\hbox{and}\qquad T^{-}\!=\,2\log R^{-}\!=\,\log\frac{|\partial M^{-}|}{4\pi}\,.$
These quantities will be instrumental in conveniently renormalizing the initial time of the IMCF that originates from $\partial M^{-}$ and for which $\partial M^{+}$ acts as an outer obstacle. Let now $w$ denote the unique solution of the problem
[…]
in the weak sense described by Xu in [74, Summary of Definitions/Theorem 1.5], where $\nu(y_{0})$ is the outward unit normal at $y_{0}\in\partial M^{+}$. As in the polar case, we set
$T^{*}\,=\,\max_{\partial M^{+}}\,w\qquad\hbox{and}\qquad T_{*}\,=\,\min_{\partial M^{+}}\,w\,$, (1.14)
and, correspondingly, $R^{*}=e^{T^{*}/2}$ and $R_{*}=e^{T_{*}/2}$. At this stage, mimiking the definition of $T^{+}_{x}$ in the connected boundary case, we define the time $T^{+}=T^{+}(M,g)$ and the associated radius $R^{+}=R^{+}(M,g)$ by
$T^{+}=\min\,\{T^{+}_{\Lambda}(p_{\Lambda}(R^{-})),T^{*}\}\qquad\hbox{and}\qquad R^{+}=\min\,\{R^{+}_{\Lambda}(p_{\Lambda}(R^{-})),R^{*}\}\,$,
and we note that they satisfy the relation $T^{+}=2\log R^{+}$.
Remark 1.8.
We first observe that, in order for $T^{+}$ and $R^{+}$ to be well defined, the argument of $p_{\Lambda}(\cdot)$ must lie in the interval $0,\sqrt{3/\Lambda}]$. Furthermore, for the interpretation of $\partial M^{-}$ as a black-hole horizon to be genuinely compelling from a physical and geometric standpoint, it is in fact preferable that this argument belongs to the more restrictive interval $0,\sqrt{1/\Lambda}]$. We note that this would be the case, under the hypotheses that will appear in Theorem 1.9. Indeed, strict stability of $\partial M^{-}$ implies $|\partial M^{-}|<4\pi/\Lambda$, or equivalently
$0<R^{-}<\sqrt{{1}/{\Lambda}}\qquad\text{and}\qquad-\infty<T^{-}<-\log\Lambda\,.$
Therefore $p_{\Lambda}(R^{-})<1/(3\sqrt{\Lambda})$ and the outer root $R_{\Lambda}^{+}(p_{\Lambda}(R^{-}))$ is well defined and strictly larger than $\sqrt{1/\Lambda}$.
Having introduced all the notation needed, we can proceed in analogy with (1.11) and define the $1$-harmonic mass of the Riemannian band $(M,g)$ as
$\mathfrak{m}_{\Lambda}^{(1)}(M,g)\,:=\,\mathfrak{m}_{\rm Haw}(\partial M^{-})\,+\,\int_{T^{-}}^{T^{+}}\left(\frac{e^{\tau/2}}{16\pi}\right)\big(4\pi-\Lambda\,{\rm Per}(\Omega_{\tau})\big)\,d\tau,$ (1.15)
where $\Omega_{\tau}=\{w<\tau\}$. We observe that this invariant can be realized as the limit, as $p\to 1$, of the $p$-harmonic masses associated with the solutions of the Dirichlet problems
$\begin{dcases}\Delta_{p}u\,=\,0&\text{in }M,\\ u\,=\,e^{-T^{-}/(p-1)}&\text{on }\partial M^{-},\\ u\,=\,0&\text{on }\partial M^{+},\end{dcases}$
for $1<p<3$, via a formal limiting procedure analogous to that described in Subsection 1.5. Furthermore, one readily verifies that the $1$-harmonic Mass of the model solutions coincides with the parameter $m$ appearing in formula (1.12). Finally, when the boundary portion $\partial M^{-}$ collapses to a point $x$, both its area and its Hawking mass converge to zero, whereas $T^{-}$ and $T_{\Lambda}^{+}(p_{\Lambda}(R^{-}))$ converge to -$\infty$ and $T_{\Lambda}=\log(3/\Lambda)$, respectively. This shows that Definition (1.15) is consistent with the one of the Polarized $1$-harmonic Mass introduced in (1.11). With this new notion at hand, we are now in a position to state the following theorem.
Theorem 1.9 (Riemannian Penrose Inequality for the $1$-harmonic Mass).
Let $(M,g)$ be a compact, three-dimensional Riemannian band, with scalar curvature satisfying
${\mathrm{R}}\geq 2\Lambda,$
for some $\Lambda>0$, and with compact minimal boundary $\partial M$ decomposing into two connected components
$\partial M\,=\,\partial M^{-}\sqcup\partial M^{+},$
such that $|\partial M^{-}|<|\partial M^{+}|$. Assume that $\partial M^{-}$ is strictly stable, that $\partial M^{+}$ is simply connected and unstable and that there are no closed minimal surfaces in $M\setminus\partial M$. Then, the $1$-harmonic Mass of $(M,g)$ defined in (1.15) satisfies
$\mathfrak{m}_{\Lambda}^{(1)}(M,g)\,\geq\,\sqrt{\frac{|\partial M^{-}|}{16\pi}}\left(1-\frac{\Lambda}{12\pi}|\partial M^{-}|\right),$ (1.16)
with equality if and only if $(M,g)$ is isometric to the Schwarzschild–de Sitter solution with mass parameter $m=\mathfrak{m}_{\Lambda}^{(1)}(M,g)$.
The inequality (1.16) is formally compatible with the classical Riemannian Penrose inequality for time-symmetric, asymptotically flat initial data with vanishing cosmological constant $\Lambda=0$. It is likewise consistent with the Penrose-type inequalities formulated by Bray and Chruściel in [18] for time-symmetric, asymptotically hyperbolic initial data with negative cosmological constant. In the case where the outermost minimal boundary is connected, these inequalities take the form
$\mathfrak{m}_{\rm Haw}(M,g)\,\geq\,\sqrt{\frac{|\partial M|}{16\pi}}\left(\frac{1}{2}\,\chi(\partial M)\;-\;\frac{\Lambda}{12\pi}\,|\partial M|\right)\,$,
where $\chi(\partial M)$ denotes the Euler characteristic of the boundary and $\mathfrak{m}_{\rm Haw}(M,g)$ denotes the total Hawking mass of the manifold. As already discussed, this concept is unfit to provide a reliable concept of mass when $\Lambda>0$. The above theorem suggest that the $1$-harmonic mass could provide a valid substitute.
@@
In Section 2,we establish monotonicity formulas on Riemannian bands under the assumption of a lower bound on the scalar curvature. These formulas are substantially more general than those required for the arguments developed in the rest of the paper. Nonetheless, they serve a dual purpose: on the one hand, they situate our contribution within a broader research perspective; on the other hand, they are expected to play a central role in forthcoming extensions of the theory in several natural directions, including a comprehensive analysis of Penrose-type inequalities in the presence of a nonzero cosmological constant. In Section 3, we specialize to the case of globally defined $p$-Green’s functions with $1<p<3$. Starting from the study of rotationally symmetric solutions, we then identify a suitable choice of the structural coefficients that appear in the general formulation of the monotonicity formulas, in such a way as to characterize static zero-mass metrics (space forms). These metrics constitute the natural reference configurations for any proposed notion of mass and are therefore required to arise as the rigidity case in inequalities intended to yield positive mass theorems. In Section 4, we begin our detailed analysis of the case of a positive cosmological constant, which is the main focus of the present work, and we refine the description of the structural coefficients. These will be crucial for evaluating the limits of the monotone quantities at both the pole of the Green’s function and the boundary of the manifold. This, in turn, prepares the ground for a first group of main results, namely Theorems 1.3 and 1.4, corresponding, respectively, to Theorem 5.4 in Section 5 and Theorem 6.2 in Section 6. In Section 4, we also provide additional remarks on the behavior of the structural coefficients in the limit as $p\to 1^{+}$, which support some of the more speculative considerations presented in Subsection 1.4. These considerations underlie the notion of $1$-harmonic mass, introduced in formula (1.11) and discussed in Subsection 1.5. In this subsection we state a positive mass theorem and a Penrose inequality, both including a rigidity statement. These results are then proved in Section 7 and constitute the second group of main results of the present work. Appendix A contains auxiliary material needed for the treatment of the case $\Lambda>0$ in Theorem 3.10, as well as for the analysis carried out in Section 4. The material in Appendix B underpins the arguments developed in Sections 5 and 6.
@@ 2. Monotonicity formulas for $p$-harmonic functions on Riemannian bands
where the $p$-Laplacian is defined as $\Delta_{p}\phi={\rm div}(|\nabla\phi|^{p-2}\nabla\phi)$. The fact that the above problem admits a unique solution $u\in\mathcal{W}^{1,p}(\Omega)\cap\mathscr{C}^{0}(\overline{\Omega})$ is standard, see for example by [50, Theorem 2.16] for a proof based on the direct method in the calculus of variations. Following [56], we set $w=-(p-1)\log u$, so that $w$ satisfies
@@ 3.2. Green’s functions for the $p$-Laplacian on 3-dimensional spaceforms and the selection of $\alpha$
one concludes that $\Psi(r)>0$ for all $0<r<R_{\Lambda}$. Having all these elements at our disposal, we can finally invert relations (3.36), thereby constructing two solutions $\lambda$ and $\mu$ of system (2.4), defined for all times and exhibiting the prescribed asymptotic expansions as $t\to$-$\infty$. This ensures that these globally defined solutions are indeed the maximal extensions fo the ones constructed in Theorem 3.4 and Theorem 3.7. ∎
@@ 5. A Positive Mass Theorem for the Polarized $p$-harmonic Mass
Let $(M,g)$ be a compact three-dimensional Riemannian manifold with smooth boundary $\partial M$. For $1<p<3$, the Polarized $p$-harmonic Mass of $(M,g)$, with pole at $x\in M\setminus\partial M$, is given by
@@
Let $(M,g)$ be a compact three-dimensional Riemannian manifold with smooth connected boundary $\partial M$, whose scalar curvature $\mathrm{R}$} satisfies
@@ 6. A Positive Mass Theorem for the $p$-harmonic Total Mass
Let $(M,g)$ be a compact three-dimensional Riemannian manifold with smooth connected minimal boundary $\partial M$, whose scalar curvature $\mathrm{R}$} satisfies
@@
Theorem B.2, coupled with the Comparison Principle for $p$-harmonic functions (see, e.g.  [61, Theorem 3.5.1]), implies that for every $n$ there exists $M_{n}$ such that
@@
Fix a point $x_{0}\in\partial M$, and let $\nu$ be the unit normal vector to $\partial M$ at $x_{0}$, pointing toward the interior of $M$. It is evident that establishing the desired statement is equivalent to proving that
@@
Now, the upper bound in (6.10) coupled with the Comparison Principle for $p$-harmonic functions (see, e.g.  [61, Theorem 3.5.1]) implies that the functions $v_{\varepsilon}$ are uniformly bounded in the complement of suitably small coordinate balls centered at $x_{\varepsilon}$, namely